OpenStudy (wwe123):

the sum of the first n terms of the A.P 3,5 1/2,8....is equal to the 2nd term of the AP 16 1/2, 28 1/2 40 1/2 . calculate the value of n

OpenStudy (wwe123):

\[here 3,5\frac{ 1 }{ 2 } , 8 \] & \[16\frac{ 1 }{ 2 }, 28 \frac{ 1 }{ 2 }, 40\frac{ 1 }{ 2}\]

hartnn (hartnn):

is the question correct ? = '2nd term of AP.....' oe its '2nd n terms of AP' ??

OpenStudy (wwe123):